OpenStudy (anonymous):

1 2/5 X 4 2/3 Multiply and simplify please

OpenStudy (anonymous):

simply multiply 12/5 * 42/3 = 12 * 42 ---------- 5 * 3 (divide 12 by 3) 4*42 = ----- 5 168/5

OpenStudy (anonymous):

Are those mixed numbers or improper fractions

OpenStudy (anonymous):

7/5 x 14/3 = 98/15= 6. 8/15